# CSDO: Enhancing Efficiency and Success in Large-Scale Multi-Vehicle Trajectory Planning

### v1

@@ Abstract
This paper presents an efficient algorithm, naming Centralized Searching and Decentralized Optimization (CSDO), to find feasible solution for large-scale Multi-Vehicle Trajectory Planning (MVTP) problem. Due to the intractable growth of non-convex constraints with the number of agents, exploring various homotopy classes that imply different convex domains, is crucial for finding a feasible solution. However, existing methods struggle to explore various homotopy classes efficiently due to combining it with time-consuming precise trajectory solution finding. CSDO, addresses this limitation by separating them into different levels and integrating an efficient Multi-Agent Path Finding (MAPF) algorithm to search homotopy classes. It first searches for a coarse initial guess using a large search step, identifying a specific homotopy class. Subsequent decentralized Quadratic Programming (QP) refinement processes this guess, resolving minor collisions efficiently. Experimental results demonstrate that CSDO outperforms existing MVTP algorithms in large-scale, high-density scenarios, achieving up to 95% success rate in 50m $\times$ 50m random scenarios around one second. Source codes are released in https://github.com/YangSVM/CSDOTrajectoryPlanning.
@@ I Introduction
Multi-vehicle trajectory planning (MVTP) seeks to generate a set of collision-free trajectories for multiple vehicles, from current positions to pre-set goals in a known unstructured environment, while minimizing travel time [1, 2]. It is a fundamental problem with diverse applications, such as cooperative parking and warehouse automation. In practical applications, there is a need to efficiently obtain solutions with a limited time [3]. As a non-convex optimization problem, MVTP necessitates a trade-off between solution quality and computational efficiency [4]. Particularly, in scenarios involving a large number of vehicles within a narrow space, the frequency of vehicle-to-obstacle and vehicle-to-vehicle conflicts increases, complicating the search for optimal or even feasible solutions [5, 6]. This work aims to develop an efficient algorithm that quickly finds feasible solutions with a high success rate for large-scale MVTP problems.
@@ I-A Related Work
The homotopy class can be loosely defined as a set of solutions that are capable of continuous deformation into one another, without intersecting obstacles or other agents [7]. Exploring various homotopy classes is essential for solving large-scale MVTP problems [1]. Therefore, we will evaluate the scalability of current MVTP algorithms in narrow space with their ability to explore homotopy classes. Table I summarizes the existing MVTP algorithms.
@@
Methods Feature Runtime
Narrow Space
Scalability
@@
Distributed planning Lack of cooperation. Fast Small
@@
Coupled planning methods [8, 9] treat all vehicles as a single, very high dimensional agent. This approach relies solely on the optimizer’s capability to traverse between different homotopy classes. While coupled planning methods guarantee completeness and even optimality, the computational complexity increases rapidly with growing number of non-convex constraints. In general, coupled planning methods exhibit poor scalability.
Distributed Planning methods address MVTP in a single-agent manner, treating others as moving obstacles [10, 11, 12]. They have high efficiency in sparse scenarios but struggle with coordination, restricting the exploration of homotopy classes. In practice, these methods often struggle to generate high-quality collaboration and the success rate decreases as scale increases, particularly in obstacle-dense scenarios.
@@
Tube construction methods [15, 16, 17] construct a safe corridor for each vehicle so the vehicle can be separated with the obstacles and other vehicles. Tube construction’s solution is strictly homotopic with the reference trajectories. Therefore, it only searches limited homotopy classes and performs poorly without an approximately feasible initial guess.
Grid search [18, 2] based methods discretize vehicle poses, actions, and space, utilizing a search algorithm to find discrete trajectories. This search algorithm is closely linked to a well-studied problem known as Multi-Agent Path Finding (MAPF), focusing on planning collision-free paths for multiple agents in a grid-like environment while minimizing travel time. Despite the NP-hard nature of MAPF, various efficient sub-optimal algorithms can generate paths for hundreds of agents in under a second [19], aligning with MVTP’s need for finding feasible solutions efficiently. However, the potential of these efficient MAPF algorithms remains largely unexplored in the MVTP field [20]. Moreover, akin to single-agent grid searching motion planning algorithms, when the search step size is too small, the search space becomes too large, posing challenges for real-time requirements. Conversely, when the search step size is too large, the solution space diminishes, making it challenging to find a solution, and collisions may occur between search steps, rendering the solution infeasible. Therefore, grid search-based methods are more suitable for generating a coarse initial guess containing homotopy class information rather than directly generating fine solutions.
I-B Motivation and Contribution
Existing methods search limited homotopy classes or search them inefficiently. In this letter, we propose a hierarchical planner, centralized searching and decentralized optimization (CSDO), tailored for large-scale MVTP problem. Specifically, by adapting the efficient MAPF algorithm priority based search (PBS) [19] for MVTP to generate an initial guess, CSDO aims at searching various homotopy classes with a large grid size, despite initial minor collisions. We further employ a tube construction method to generate a safe and precise solution. By transforming multi-vehicle constraints into single-vehicle planning problems near this initial guess, we simplify the MVTP into parallel, solvable ones, enhancing efficiency.
@@ I-B Motivation and Contribution
    A seamless adaptation of the priority-based search method from the MAPF domain into the complex non-holonomic MVTP problems, enabling efficient exploration for feasible or near-feasible solutions.
@@
    An efficient distributed local solver is introduced. Given a homotopically correct reference solution, the local solver can generate feasible solutions within approximately 300 ms for nearly 100 agents on a typical 100 m × 100 m size random map.
@@ II Problem Definition
A MVTP problem can be defined by a ten element tuple $\langle M,\mathcal{W},\mathcal{O},z,\mathcal{R},s,g,f,\mathcal{T},P\rangle$. Consider a system consisting of $M$ front-steering agents $a^{(1)},a^{(2)},...,a^{(M)}$ operating in a continuous planar workspace $\mathcal{W}\subset\mathbb{R}^{2}$. We use $[M]$ to denote the set $\{1,2,...,M\}$ for simplicity. There is some random static obstacles lying in the environment and occupying the workspace $\mathcal{O}$. We use superscript $(i)$ to represent the variable related to agent $a^{(i)}$. $z^{(i)}=[x^{(i)},y^{(i)},\theta^{(i)}]^{T}\in\mathbb{R}^{3}$ refers to the state of agent $a^{(i)}$, where $(x^{(i)},y^{(i)})$ is the position of rear axis center of agent $a^{(i)}$ and $\theta^{(i)}$ is its yaw angle. We use agent occupancy function $\mathcal{R}(z):\mathbb{R}^{3}\rightarrow\mathcal{W}$ to represent the workspace occupied by the agent’s body at state $z$. Agent’s trajectory is represented by a sequence of vehicle states sampled at fixed time interval $\Delta t$. It is denoted as $\mathcal{T}=[z_{0},z_{1},...,z_{\tau}]$, where the $\tau+1$ is the number of states in the trajectory. For one MVTP task, agent $a^{(i)}$’s trajectory $\mathcal{T}^{(i)}$ need to start from the start state $s^{(i)}$ and end at the goal state $g^{(i)}$. The MVTP task is finished only when all the agents arrive at their goals at a specific time $\tau_{f}$.
$z^{(i)}_{0}=s^{(i)},z^{(i)}_{\tau_{f}}=g^{(i)},\forall i\in[M].$ (1)
The control input of agent $a^{(i)}$ is denoted as $u^{(i)}=[v^{(i)},\omega^{(i)}]^{T}$, where $\omega^{(i)}=\dot{\phi}^{(i)}$, $\phi^{(i)}$ is the front-wheel steering angle, $v^{(i)}$ is the velocity. For agent $a^{(i)}$, its planned trajectory $z^{(i)}_{t}$ should be kinematic feasible for the Ackermann-steering model $f$. We denote $\zeta=[x,y,\theta,\phi]^{T}$ to include the state and steering angle. The kinematic model can be written as follows,
$\zeta^{(i)}_{t+1}$ $=\zeta^{(i)}_{t}+\begin{bmatrix}v^{(i)}_{t}\cos(\theta^{(i)}_{t})\\ v^{(i)}_{t}\sin(\theta^{(i)}_{t})\\ v^{(i)}_{t}\tan(\phi^{(i)}_{t})/L\\ \omega\\ \end{bmatrix}\Delta t,$ (2)
$\forall 0$ $\leq t<\tau_{f},i\in[M],$
$\left|v^{(i)}_{t}\right|\leq v_{max},\left|\omega^{(i)}_{t}\right|\leq\omega_{max},\forall 0\leq t<\tau_{f},i\in[M],$ (3)
$\left|\phi^{(i)}_{t}\right|\leq\phi_{max},\forall 0\leq t\leq\tau_{f},i\in[M],$ (4)
where the $L$ is the vehicle wheelbase. The agents cannot collide with static obstacles or any other agents, i.e.
@@
The solution plan $P$, comprising collision-free trajectories for all agents, is measured by its $makespan$ $\tau_{f}$. Considering the summarized elements, a traditional optimal control problem can be formulated as
Minimize $\quad\quad\tau_{f}$ (7)
@@ III Method
The overall CSDO framework is illustrated in Fig. 1. Upon receiving the start poses, goal poses, and obstacle information, these components are combined to form a Multi-Vehicle Trajectory Planning (MVTP) instance. Subsequently, the centralized priority based searching phase generates coarse trajectories as an initial guess. The decentralized Quadratic Programming (QP) refinement follows, where inter-vehicle constraints are divided and solved in parallel using the QP formulation. Finally, the obtained solution is then sent to the respective multiple vehicles, which subsequently track the collision-free trajectories to reach their goals.
@@ III-A Centralized Priority based Searching
By discretizing the MVTP problem, we can utilize search algorithms to find an initial guess. As in Fig. 2, the deiscretizing process consists of state $z$ discretization, kinematic model $f$ discretization and agent collision detection implementation in formula (5-6) as in single agent search algorithm hybrid A* [21]. The state is still stored as a continuous vector but only a limited number of states are kept. As in Fig. 2(a), we discretize the workspace $\mathcal{W}$ into $n_{xy}$ grids. Every grid only has $n_{\theta}$ angles. Then at most $n_{xy}\times n_{\theta}\times(\tau_{f}+1)\times M$ states can be stored. For every agent at a specific time, its continuous state $z$ will calculate a corresponding discrete state $\hat{z}=[\hat{x},\hat{y},\hat{\theta}]^{T}$ based on the L2 loss, where $(\hat{x},\hat{y})$ is the center position of the grid and $\hat{\theta}\in\{0,2\pi/n_{\theta},...,2\pi(n_{\theta}-1)/n_{\theta}\}$. As in Fig. 2(a), if two states $z_{1}^{(i)}(t_{k})$, $z_{2}^{(i)}(t_{k})$ of same agent and timestamp have the same discrete state $\hat{z^{(i)}(t_{k})}$, then only the state with minimum cost is saved. The kinematic model $f$ is simplified to an action set. $Actions=\{FL,FS,FR,BL,BS,BR,Wait\}$, which stand for front-max-steering-left, front-straight, front-max-steering-right, back-max-steering-left, back-straight, back-max-steering-right and wait respectively. Except for the $Wait$ action, all the actions, travels for the same distance. Collision detection utilizes the Separating Axis Theorem and checks only if vehicles collide at the sampling moment, without considering collisions between state transitions.
@@
Our overall framework is illustrated in Fig. 3. Centralized searching is divided into two layers of search. In the high level, each node represents a subproblem, and each node contains a set of partial priority orders as constraints. Specifically, for any two agents, if $i\prec j$, then $j$ has a lower priority than $i$, and $j$ treats $i$ as a dynamic obstacle to avoid in the low-level planner. For any two agents, they may not have a relationship of $i\prec j$ or $j\prec i$, meaning they ignore each other and may have collisions. If collisions arise within a node’s plan, two possible partial orders according to a specific conflict will be added to generate child nodes until collisions are entirely resolved.
@@ 4 STACK $\leftarrow\{Root\}$;
9      $(i,j)\leftarrow$ one collision in $N.plan$;
@@ 10       $N_{i},N_{j}\leftarrow N$ ;
13             $\boldsymbol{\prec}_{N_{x}}\leftarrow\boldsymbol{\prec}_{N}\cup\{y\prec x\}$ ;
@@ 14             UPDATEPLAN($N_{x}$);
15
16      Insert $N_{i}$ and $N_{j}$ into STACK in descending order of their makespan;
17
@@ 16      Insert $N_{i}$ and $N_{j}$ into STACK in descending order of their makespan;
We adapt PBS [19, 22] to address our problem as follows. As in algorithm 1, the root node [Line 1] is initialized with an empty set of priority orderings, but we employ a random restart technique to speed up the search process. Within the root node, we attempt to plan the agents sequentially, treating the previously planned agents as dynamic obstacles. If any agent encounters planning failure due to obstruction by preceding agents, we allow it to plan its trajectory freely; When expanding a node, we check for collisions between each vehicle pair at every time step. If no collisions are detected, the node is considered the final result [Line 5-6]. Otherwise, we select a pair of colliding vehicles, $i$ and $j$ [Line 7]. Two constraints, $i\prec j$ and $j\prec i$, are created and resulting in two child nodes. A low-level planner is then employed to replan the trajectories to satisfy the new priorities. To avoid redundant replanning, we topologically sort the new priorities in descending order and plan conflicting lower-priority vehicles from higher to lower priority [Line 11]. Finally, child nodes are inserted [Line 12] in non-decreasing order of the planned makespan of the nodes.

### v2

@@ Abstract
This paper presents an efficient algorithm, naming Centralized Searching and Decentralized Optimization (CSDO), to find feasible solution for large-scale Multi-Vehicle Trajectory Planning (MVTP) problem. Due to the intractable growth of non-convex constraints with the number of agents, exploring various homotopy classes that imply different convex domains, is crucial for finding a feasible solution. However, existing methods struggle to explore various homotopy classes efficiently due to combining it with time-consuming, precise trajectory solution finding. CSDO, addresses this limitation by separating them into different levels and integrating an efficient Multi-Agent Path Finding (MAPF) algorithm to search homotopy classes. It first searches for a coarse initial guess using a large search step, identifying a specific homotopy class. Subsequent decentralized Sequential Quadratic Programming (SQP) refinement processes this guess, resolving minor collisions efficiently. Experimental results demonstrate that CSDO outperforms existing MVTP algorithms in large-scale, high-density scenarios, achieving up to a 95% success rate in 50 m $\times$ 50 m random scenarios around one second. Source codes are released in https://github.com/YangSVM/CSDOTrajectoryPlanning.
@@ I Introduction
Multi-Vehicle Trajectory Planning (MVTP) seeks to generate a set of collision-free trajectories for multiple vehicles, from current positions to pre-set goals in a known unstructured environment, while minimizing travel time [1]. It is a fundamental problem with diverse applications, such as cooperative parking and warehouse automation. In practical applications, there is a need to efficiently obtain solutions within a limited time [2]. As a non-convex optimization problem, MVTP necessitates a trade-off between solution quality and computational efficiency [3]. Particularly in scenarios involving a large number of vehicles, the frequency of vehicle-to-obstacle and vehicle-to-vehicle conflicts increases, complicating the search for optimal or even feasible solutions [4]. This work aims to develop an efficient algorithm that quickly finds feasible solutions with a high success rate for large-scale MVTP problems.
@@ I-A Related Work
Existing works struggle to find feasible solutions quickly at large scales, as shown in Table I. The main challenge lies in efficiently exploring various homotopy classes [1]. The homotopy class can be loosely defined as a set of solutions that are capable of continuous deformation into one another, without intersecting obstacles or other agents [5]. Different homotopy classes can be seem as combinations of various routes and agent behaviors. The quality of local optimal solutions within each homotopy class varies significantly. In large-scale scenarios, only a few homotopy classes might contain feasible solutions, making the exploration of various homotopy classes essential [1]. Therefore, we evaluate the scalability of current MVTP algorithms in dense space based on their ability to explore homotopy classes.
@@
Methods Feature Runtime Scalability
@@
Distributed Lack of cooperation. Fast Small
Sampling based
Probabilistic complete;
Asymptotically optimal
Relative fast Relative large
@@
Coupled planning methods [6] treat all vehicles as a single, high dimensional agent. This approach relies solely on the optimizer’s capability to traverse between different homotopy classes. While coupled planning methods guarantee completeness and optimality, the computational complexity increases rapidly with the growing number of non-convex constraints. In general, coupled planning methods exhibit poor scalability.
Distributed Planning methods address MVTP in a single-agent manner, treating others as moving obstacles [7, 8, 9], or achieve collision avoidance through communication [10, 11]. They have high efficiency in sparse scenarios but struggle with coordination, restricting the exploration of homotopy classes. In practice, these methods often struggle to generate high-quality collaboration, and the success rate decreases as scale increases, particularly in obstacle-dense scenarios.
Sampling-based methods [12, 13, 14] mainly extend the Probabilistic Road Map (PRM) and Rapidly-exploring Random Tree (RRT) to multi-robot systems. These methods can provide probabilistic completeness and even asymptotical optimality. However, in crowded scenarios, these methods require a large number of samples, which can still lead to timeouts.
@@
Tube construction methods [17, 18, 19] construct a safe corridor for each vehicle so the vehicle can be separated from the obstacles and other vehicles. Tube construction’s solution is strictly homotopic to the reference trajectories. Therefore, it only searches limited homotopy classes and performs poorly without an approximately feasible initial guess.
Grid search [4, 3] based methods discretize vehicle poses, actions, and space, utilizing a search algorithm to find discrete trajectories. This search algorithm is closely linked to a well-studied problem known as Multi-Agent Path Finding (MAPF), focusing on planning collision-free paths for multiple agents in a grid-like environment while minimizing travel time. Despite the NP-hard nature of MAPF, various efficient sub-optimal algorithms can generate paths for hundreds of agents in under a second [20], aligning with MVTP’s need for finding feasible solutions efficiently. However, the potential of these efficient MAPF algorithms remains largely unexplored in the MVTP field [21]. Moreover, akin to single-agent grid searching motion planning algorithms, when the search step size is too small, the search space becomes too large, posing challenges for real-time requirements. Conversely, when the search step size is too large, the solution space diminishes, making it challenging to find a solution, and collisions may occur between search steps, rendering the solution infeasible. Therefore, grid search-based methods are more suitable for generating a coarse initial guess containing homotopy class information than directly generating fine solutions.
I-B Motivations and Contributions
Based on the aforementioned literature review, existing methods search limited homotopy classes or search them inefficiently. In this letter, we propose using an efficient MAPF solver to explore various coarse initial guesses with a large step size, which implictly encode different homotopy classes. After obtaining an initial guess that contains a specific homotopy class, decentralized optimization is employed to quickly generate a nearby kinematically feasible solution, ultimately achieving rapid generation of feasible solutions within a limited time.
@@ I-B Motivations and Contributions
    A seamless adaptation of the priority-based search method from the MAPF domain into the complex non-holonomic MVTP problems, enables efficient exploration for feasible or near-feasible solutions.
@@
    An efficient distributed local solver is introduced. Given a homotopically correct reference solution, the local solver can generate feasible solutions quickly.
@@ II Problem Definition
The MVTP problem can be defined by a ten element tuple $\langle M,\mathcal{W},\mathcal{O},z,\mathcal{R},s,g,f,\mathcal{T},X\rangle$. Consider a system consisting of $M$ front-steering agents $a^{(1)},a^{(2)},...,a^{(M)}$ operating in a continuous planar workspace $\mathcal{W}\subset\mathbb{R}^{2}$. For simplicity, we use $[M]$ to denote the set $\{1,2,...,M\}$ and superscript $(i)$ to represent the variable related to agent $a^{(i)}$. There are some random static obstacles lying in the environment and occupying the workspace $\mathcal{O}$. $z=[x,y,\theta,\phi]^{T}\in\mathbb{R}^{4}$ refers to the state, where $(x,y)$ is the position of rear axis center, $\theta$ is yaw angle and $\phi$ is front-wheel steering angle. The control input is denoted as $u=[v,\omega]^{T}\in\mathbb{R}^{2}$, where $\omega=\dot{\phi}$, $v$ is the velocity. Agent $a^{(i)}$’s trajectory is represented by a sequence of its states sampled at fixed time interval $\Delta t$. It is denoted as $\mathcal{T}^{(i)}=[z_{0}^{(i)},z_{1}^{(i)},...,z_{\tau_{f}^{(i)}}^{(i)}]$, where the $\tau_{f}^{(i)}+1$ is the number of states in the trajectory. For one MVTP task, $\mathcal{T}^{(i)}$ need to start from the start state $s^{(i)}$ and end at the goal state $g^{(i)}$.
$z^{(i)}_{0}=s^{(i)},z^{(i)}_{\tau_{f}^{(i)}}=g^{(i)},\forall i\in[M].$ (1)
The task finish time is $makespan$ $\tau_{f}$, where $\tau_{f}=\max_{i\in[M]}\tau_{f}^{(i)}$. It is assumed that the agent waits at the goal until all the agents have reached their goals, i.e. $z_{t}^{(i)}=g^{(i)},\forall\tau_{f}^{(i)}\leq t\leq\tau_{f}$. The planned trajectory $\mathcal{T}$ should be kinematic feasible for the Ackermann-steering model $f$, i.e.,
$z_{t+1}$ $=z_{t}+\begin{bmatrix}v_{t}\cos\theta_{t}\\ v_{t}\sin\theta_{t}\\ v_{t}\tan(\phi_{t})/L\\ \omega\\ \end{bmatrix}\Delta t,\forall 0\leq t<\tau_{f},$ (2)
$\left|v_{t}\right|\leq v_{max},\left|\omega_{t}\right|\leq\omega_{max},\forall 0\leq t<\tau_{f},$ (3)
$\left|\phi_{t}\right|\leq\phi_{max},\forall 0\leq t\leq\tau_{f},$ (4)
where the $L$ is the vehicle’s wheelbase. We use agent occupancy function $\mathcal{R}(z):\mathbb{R}^{4}\rightarrow\mathcal{W}$ to represent the workspace occupied by the agent’s body at state $z$. The agents cannot collide with static obstacles or any other agents, i.e.,
@@
The solution plan $X$ comprises collision-free trajectories and control inputs of all agents. The solution quality is measured by $\tau_{f}$. Considering the summarized elements, a traditional optimal control problem [1] can be formulated as
$\min_{X,\tau_{f}}$ $\quad\quad\tau_{f}$ (7)
@@ III Method
The overall CSDO framework is illustrated in Fig. 1. Upon receiving the start poses, goal poses, and obstacle information, these components are combined to form a Multi-Vehicle Trajectory Planning (MVTP) instance. Subsequently, the centralized priority based searching phase generates coarse trajectories as an initial guess. The decentralized Sequential Quadratic Programming (SQP) refinement follows, where inter-vehicle constraints are decomposed and sent to multiple vehicles. Each vehicle utilizes the SQP solver to derive its trajectory and start to tracking simultaneously to reach their respective goals.
@@ III-A Centralized Priority based Searching
By discretizing the MVTP problem, we can utilize search algorithms to find an initial guess. As in Fig. 2, the discretizing process consists of state $z$ discretization, kinematic model $f$ discretization and agent collision detection implementation in formula (5-6) as in single vehicle search algorithm hybrid A* [22]. The discrete state $\hat{z}=[\hat{x},\hat{y},\hat{\theta},0]^{T}$, which contains the closest grid center position and discrete yaw angle, i.e., $\hat{z}=\operatorname*{arg\,min}_{\hat{z}}||z-\hat{z}||^{2}$. As in Fig. 2(a), $\hat{z}^{(i)}(t_{k})$ is the discrete state of $z^{(i)}(t_{k})$; The kinematic model $f$ is simplified to an action set and limited to constant speed due to time complexity. $Actions=\{FL,FS,FR,BL,BS,BR,Wait\}$, which stand for front-max-steering-left, front-straight, front-max-steering-right, back-max-steering-left, back-straight, back-max-steering-right and wait respectively. Except for the $Wait$ action, all the actions, travel for the same step size $\Delta S$. In the grid search process, we adopt a large step size to search a coarse trajectory; Separating Axis Theorem is utilized the to check the collision at the sampling moment.
@@
Our overall framework is illustrated in Fig. 3. It is worth noting that our CSDO architecture can achieve optimal or bounded sub-optimal solutions by applying corresponding MAPF algorithms. We choose PBS to find a feasible solution efficiently. Centralized searching is divided into two layers of search. At the high level, each node represents a subproblem, and each node contains a constraint set consisting of partial priority orders. Specifically, a partial priority order consists of two agents and refers to the avoidance relationship. For instance, if $a^{(i)}\prec a^{(j)}$, then $a^{(j)}$ has a lower priority than $a^{(i)}$, and $a^{(j)}$ treats $a^{(i)}$ as a dynamic obstacle to avoid in the low-level planner. For any two agents, they may not have a relationship of $a^{(i)}\prec a^{(j)}$ or $a^{(j)}\prec a^{(i)}$, meaning they ignore each other and may have collisions. If collisions arise within a node’s plan, two possible partial orders according to a specific conflict will be added to generate child nodes until collisions are entirely resolved.
@@ 4 STACK $\leftarrow\{Root\}$;
9      $(a^{(i)},a^{(j)})\leftarrow$ one colliding agent pair in $n.plan$;
10       /* Node expansion. */
@@ 11       $n_{i},n_{j}\leftarrow n$ ;
14             $\boldsymbol{\prec}_{n_{x}}\leftarrow\boldsymbol{\prec}_{n}\cup\{a^{(x)}\prec a^{(y)}\}$ ;
@@ 15             UPDATEPLAN($n_{x}$);
16
17      Insert feasible nodes into STACK in descending order of their makespan;
@@
We adapt PBS [20, 23] to address our problem as follows. As in algorithm 1, the root node [Line 1] is initialized with an empty set of priority orders, but we employ a warm start technique to speed up the search process. Within the root node, we attempt to plan the agents sequentially, treating the previously planned agents as dynamic obstacles. If any agent encounters planning failure due to obstruction by preceding agents, we allow it to plan its trajectory freely; When expanding a node, we check for collisions between each vehicle pair. If no collisions are detected, the node is considered the final result [Lines 5 to 6]. Otherwise, we select a pair of colliding vehicles, $a^{(i)}$ and $a^{(j)}$ [Line 7]. The following part describes the detailed procedure of node expansion as shown in Fig. 3. Two constraints, $a^{(i)}\prec a^{(j)}$ and $a^{(j)}\prec a^{(i)}$, are created and added separately to form the new child nodes’ constraint sets. For example, $\prec_{n_{i}}=\prec_{n}\cup\{a^{(i)}\prec a^{(j)}\}$ as in node $n_{i}$ [Lines 10 to 11]. One straightforward replanning method involves replanning all the agents according to $n_{x}$. However, to update the plan [Line 12] without redundant replanning, we first identify the agents violating the new constraint set $\prec_{n_{x}}$ and perform a topological sorting on the agents. Next, we replan the agents from higher to lower priority using the low-level planner. Finally, feasible child nodes are inserted [Line 13] in the non-decreasing order of the planned makespan of the nodes.